OpenStudy (anonymous):

If you have 400.0 mL of water at 25.00 °C and add 110.0 mL of water at 95.00 °C, what is the final temperature of the mixture?

ganeshie8 (ganeshie8):

use enthalpy formula \(q = m \ c\ \Delta t\)

ganeshie8 (ganeshie8):

energy lost by hot water = energy absorbed by cool water

OpenStudy (anonymous):

How do you set up the entire equation though?

ganeshie8 (ganeshie8):

.4 * c * (T-25) = -.11 * c * (T-95)

ganeshie8 (ganeshie8):

cancel c, solve T

OpenStudy (anonymous):

Got it. Thanks so much!

ganeshie8 (ganeshie8):

yw ! remember this while doing enthalpy related problems : q system = -q surroundings

OpenStudy (anonymous):

Actually I have one more question. Why did you use .4 and .11 instead of 400 and 110?

ganeshie8 (ganeshie8):

yeah we can use 400 and 110 also here, it doesnt matter here

ganeshie8 (ganeshie8):

ideally we must convert mL to grams. becoz in enthalpy equation, m represents mass in grams.

ganeshie8 (ganeshie8):

the conversion to grams doesnt matter here since 1 mL = 1 g, for water

ganeshie8 (ganeshie8):

the correct way of putting the equation would be : 400g * c * (T-25) = 110g * c * (T-95)

OpenStudy (anonymous):

I ended up getting 19.62 degrees, but it's incorrect :/

ganeshie8 (ganeshie8):

check ur calculation again

ganeshie8 (ganeshie8):

400g * c * (T-25) = -110g * c * (T-95) 400 * (T-25) = -110* (T-95)

ganeshie8 (ganeshie8):

im getting 40.1

OpenStudy (anonymous):

That's what I got this time too. Thank you so much! I appreciate all your help.

ganeshie8 (ganeshie8):

u getting 19.62 still ? could u show ur calculation so that i can see whats happening at ur end :)

OpenStudy (anonymous):

Nope, I got the correct answer this time. 40.1 degrees Celsius

ganeshie8 (ganeshie8):

sounds great !